A particle Q of mass m falls under gravity in a medium, whose resistance to motion is of magnitude mkv^2, where v is the speed of Q and k is a positive constant. Given that the maximum speed of Q is U, show that U=√(g/k) (g is the acceleration due to gravity).
Given that Q is projected vertically upwards with speed V(>U), show that the speed of Q is equal to U when the height of Q above the point of projection is U^2/2g ln⁡[1/2 (1+V^2/U^2 )].
Find in terms of U, V and g, the time taken for the speed of Q to decrease from V to U.
1
Expert's answer
2020-06-11T10:34:25-0400

As per the given question,

The falling particle Q of mass is m, is experiencing a resistiv force "=mkv^2"

Here v is the speed and k is the positive constant

Now applying the force balance equation of the net resultant force,

"\Rightarrow F=mg- mkv^2"


"\Rightarrow m \frac{dx^2}{dt^2}=m(g-kv^2)"


"\Rightarrow \frac{dx^2}{dt^2}=(g-kv^2)"

Now we can write this as,

"\Rightarrow \frac{d}{dt}(\frac{dx}{dt})=(g-kv^2)"


"\Rightarrow" "\frac{dv}{dt}=g-kv^2"

Now applying the maxima and minima concept,

"\frac{dv}{dx}=0"

"\Rightarrow g-kv^2=0"

"\Rightarrow v=\sqrt{\frac{g}{k}}"

ii)

"\Rightarrow \frac{dv}{dt}=g-kv^2"


"\Rightarrow \frac{dv}{g-kv^2}=dt"

now after the integration,

"t=\\frac{U^2}{2g} \\ln\u2061[\\frac{1}{2 (1+V^2\/U^2 )}]."